what is the expect value if the annual budget is $134,685,340 and the percentage spent is 8.33% and what is the variance if the actual for first month is $11,000,000?
1
Expert's answer
2013-06-21T08:02:01-0400
The expected value will be 134,685,340*0.0833 =$11,219,289.
The variance will be equal: (0.0833 -11,000,000/134,685,340)^2 =
(0.0833 - 0.0817)^2 = 0.016^2 = 0.000256 = 0.0256% So, we can state, that the variance is low and the expected value was calculated close
enough to the real value.